# SLMRec: Empowering Small Language Models for Sequential Recommendation

### v1

@@ 7 Conclusions and Outlooks
This paper explores the effectiveness of large language models in sequential recommendation. Motivated by empirical insights, we adopt vanilla knowledge distillation methods to improve the performance of small language models. With only 13% parameters and yields 8x acceleration against the LLMRec baseline, our SLMRec achieves slightly better performance. On top of our technical contributions, we believe the results in this paper could shed light on a new promising direction for building effective and efficient recommenders based on LLMs, which is largely under-explored.
@@
Appendix A Appendix
A.1 Training Details
@@ A.2 Compared methods
GRU4Rec [18] tackles the issue of modeling sparse sequential data while also adapting RNN models to the recommender system. To achieve this, the authors propose a new ranking loss function that is specifically designed for training these models. The implementation of GRU4Rec in PyTorch can be found at the URL  [^6] https://github.com/hungpthanh/GRU4REC-pytorch.
@@
BERT4Rec [46] designs a bidirectional self-attention network to model user behavior sequences. To prevent information leakage and optimize the training of the bidirectional model, a Cloze objective is used to predict the randomly masked items in the sequence by considering both their left and right context. The implementation of BERT4Rec in PyTorch can be found at the URL  [^7] https://github.com/jaywonchung/BERT4Rec-VAE-Pytorch.
@@
SASRec [29] is a self-attention based sequential model that addresses the challenge of balancing model parsimony and complexity in recommendation systems. By using an attention mechanism, SASRec identifies relevant items in a user’s action history and predicts the next item based on relatively few actions, while also capturing long-term semantics like an RNN. This enables SASRec to perform well in both extremely sparse and denser datasets. The implementation of SASRec in PyTorch can be found at the URL  [^8] https://github.com/pmixer/SASRec.pytorch.
@@
Open-P5 [54] is an open-source platform introduced to catalyze research in LLM-based generative recommender systems. It supports key model architectures like T5 and Llama-2 across diverse public datasets, focusing on sequential and straightforward recommendation tasks. The platform emphasizes the role of item IDs through various indexing methods and offers a customizable, efficient, and standardized environment for developing and assessing recommender systems. The implementation of Open-P5 in PyTorch can be found at the URL   [^9] https://github.com/agiresearch/OpenP5.
E4SRec [35] integrate of Large Language Models (LLMs) into sequential recommendation systems, offering a significant leap in handling item IDs and personalization. In the original paper, they use Softmax layer to output each user-item prediction score. The implementation of E4SRec in PyTorch can be found at the URL  [^10] https://github.com/HestiaSky/E4SRec.
A.3 Model Efficiency

### v3

@@ 8 Conclusions
This paper explores the effectiveness of large language models (LLMs) in sequential recommendation. Our motivational experiments reveal that intermediate layers in LLMs are redundant for achieving optimal recommendation performance. Motivated by empirical insights, we adopt vanilla knowledge distillation methods to improve the performance of small language models. Achieving only 13% of the parameters compared to the LLMRec baseline, our SLMRec model yields an 8x acceleration and slightly better performance. On top of our technical contributions, we believe the results in this paper could shed light on a new promising direction for building effective and efficient recommenders based on LLMs, which is largely under-explored. Additionally, we provide theoretical justifications showing that while multi-layer models optimize distinct objectives at each layer, this can introduce redundancy compared to a single-layer model that achieves its objective in one step. These theoretical insights align with the motivation behind our framework design, where we employ knowledge distillation (KD) to guide the one-layer network, enabling it to streamline the learning process and replicate the feature extraction capabilities of a multi-layer network.
@@
Appendix A Proof
A.1 Proof for Proposition 1
Proposition 1. Given the matrix $\hat{\mathbf{A}}^{(k)}=\mathbf{A}^{(k)}+\mathbf{I}$, Eqn. 7 is equivalent to a gradient descent step with step size 1 for the following optimization problem:
$\min_{H}\left\|\mathbf{H}-\hat{\mathbf{A}}^{(k)}\mathbf{H}^{(k-1)}\right\|_{2}^{2}$ (10)
Proof.
The cost function at the $k$-th layer is denoted by
$E(\mathbf{H};\mathbf{H}^{(}k-1))=\min_{\mathbf{H}}\left\|\mathbf{H}-\hat{\mathbf{A}}^{(k)}\mathbf{H}^{(k-1)}\right\|_{2}^{2}$ (11)
Then, the gradient of $E(\mathbf{H};\mathbf{H}^{(}k-1))$ is computed as
$\frac{\partial E(\mathbf{H};\mathbf{H}^{(}k-1))}{\partial\mathbf{H}}=2\left(\mathbf{H}-\hat{\mathbf{A}}^{(k)}\mathbf{H}^{(k-1)}\right)$ (12)
With the step size 1 of the gradient descent, it minimizes the cost function $E(\mathbf{H};\mathbf{H}^{(}k-1))$ at the current layer is
$\mathbf{H}^{(k)}$ $=\mathbf{H}^{(k-1)}-\frac{\partial E(\mathbf{H};\mathbf{H}^{(}k-1))}{\partial\mathbf{H}}|_{\mathbf{H}=\mathbf{H}^{(k-1)}}$ (13)
$=\mathbf{H}^{(k-1)}-\left(\mathbf{H}^{(k-1)}-\hat{\mathbf{A}}^{(k)}\mathbf{H}^{(k-1)}\right)$ (14)
$=\hat{\mathbf{A}}^{(k)}\mathbf{H}^{(k-1)}$ (15)
$=\mathbf{H}^{(k-1)}+\mathbf{A}^{(k)}\mathbf{H}^{(k-1)}$ (16)
A.2 Proof for Proposition 2
Proposition 2. For any $K$-layer attention model (where $K$ is an arbitrary positive integer) with the layer-wise updating rule defined by Eqn. 7, there exists $\mathbf{C}^{*}$ such that one gradient descent step for the optimization problem (from the initial embeddings $\mathbf{H}^{(0)}$)
$\min_{\mathbf{H}}\left\|\mathbf{H}-\mathbf{C}^{*}\mathbf{H}^{(0)}\right\|_{2}^{2},$ (17)
where $\mathbf{C}^{*}$ associated with $\mathbf{A}$, can yield the output embeddings $\mathbf{H}^{(K)}$ of the $K$-layer model.
Proof.
Similar to Theorem 1, we define $\hat{\mathbf{A}}^{(k)}$ to simpify the Eqn. 7.
$\hat{\mathbf{A}}^{(k)}=\mathbf{I}+\mathbf{A}^{(k)},$ (18)
Then Eqn. 7 can be equivalently written as
$\mathbf{H}^{(k)}=\hat{\mathbf{A}}^{(k)}\mathbf{H}^{(k-1)},$ (19)
By stacking $K$ layers of propagation, we can denote the output embeddings as
$\mathbf{H}^{(K)}=\hat{\mathbf{A}}^{(K)}\mathbf{H}^{(K-1)}=\hat{\mathbf{A}}^{(K)}\hat{\mathbf{A}}^{(K-1)}\mathbf{H}^{(K-2)}=\cdots=\hat{\mathbf{A}}^{(K)}\cdots\hat{\mathbf{A}}^{(1)}\mathbf{H}^{(0)}=\mathbf{A}^{*}\mathbf{H}^{(0)},$ (20)
where $\mathbf{A}^{*}$ defined as multiple matrix production.
We can show that solving the denoising problem with gradient step size $\frac{\mu^{*}}{2}$ w.r.t. the objective
$\min_{\mathbf{H}}\left\|\mathbf{H}-\mathbf{C}^{*}\mathbf{H}^{(0)}\right\|_{2}^{2},$ (21)
Defining $\mathbf{C}^{*}=\frac{1}{\mu^{*}}\left(\mathbf{A}^{*}-(1-\mu^{*})\mathbf{I}\right)$, $\mathbf{H}^{(k)}=\mathbf{A}^{*}\mathbf{H}^{(0)}$ will induce the output embeddings $\mathbf{H}^{(K)}$, by noticing that
$\mathbf{H}^{(K)}=$ $\mathbf{H}^{(0)}-\frac{\mu^{*}}{2}\left.\frac{\partial E(\mathbf{H};\mathbf{H}^{(0)})}{\partial\mathbf{H}}\right|_{\mathbf{H}=\mathbf{H}^{(0)}}$ (22)
$=$ $\mathbf{H}^{(0)}-2\frac{\mu^{*}}{2}(\mathbf{H}^{(0)}-\mathbf{C}^{*}\mathbf{H}^{(0)})$ (23)
$=$ $\mathbf{H}^{(0)}-2\frac{\mu^{*}}{2}[\mathbf{H}^{(0)}-\frac{1}{\mu^{*}}\left(\mathbf{A}^{*}-(1-\mu^{*})\mathbf{I}\right)\mathbf{H}^{(0)}]$ (24)
$=$ $\mathbf{H}^{(0)}-\mu^{*}\mathbf{H}^{(0)}+\mathbf{A}^{*}\mathbf{H}^{(0)}-\mathbf{H}^{(0)}+\mu^{*}\mathbf{H}^{(0)}$ (25)
$=$ $\mathbf{A}^{*}\mathbf{H}^{(0)}$ (26)
Appendix B Experiments
B.1 Motivation Experiment Results
(a) Movie (Infer) - HR@10
(b) Movie (Infer) - NDCG@10
(c) Movie (Infer) - MRR
(d) Movie (Train) - HR@10
(e) Movie (Train) - NDCG@10
(f) Movie (Train) - MRR
Figure 5: We present the relationship between the number of decoder layers and the final recommendation performance, with the performance of SASRec plotted as a baseline. Figures (a)-(c) show the results of directly using representations from the middle layers for inference without training, while (d)-(f) prune the later layers and train a model using only the specified number of layers. From the results, we observe that deeper decoder layers introduce redundancy in recommendation tasks, with models utilizing fewer layers (8-layer) achieving performance nearly equivalent to (24-layer) models.
B.2 Training Details
@@ B.3 Compared methods
Caser (Tang & Wang, 2018) introduces a novel approach to sequential recommendation systems by modeling user-item interactions as sequences, which is designed to predict the next item a user may interact with by capturing both short-term and long-term dependencies in user behavior.
@@
GRU4Rec (Hidasi et al., 2015) tackles the issue of modeling sparse sequential data while also adapting RNN models to the recommender system. To achieve this, the authors propose a new ranking loss function that is specifically designed for training these models. The implementation of GRU4Rec in PyTorch can be found at the URL  [^7] https://github.com/hungpthanh/GRU4REC-pytorch.
@@
BERT4Rec (Sun et al., 2019) designs a bidirectional self-attention network to model user behavior sequences. To prevent information leakage and optimize the training of the bidirectional model, a Cloze objective is used to predict the randomly masked items in the sequence by considering both their left and right context. The implementation of BERT4Rec in PyTorch can be found at the URL  [^8] https://github.com/jaywonchung/BERT4Rec-VAE-Pytorch.
@@
SASRec (Kang & McAuley, 2018) is a self-attention based sequential model that addresses the challenge of balancing model parsimony and complexity in recommendation systems. By using an attention mechanism, SASRec identifies relevant items in a user’s action history and predicts the next item based on relatively few actions, while also capturing long-term semantics like an RNN. This enables SASRec to perform well in both extremely sparse and denser datasets. The implementation of SASRec in PyTorch can be found at the URL  [^9] https://github.com/pmixer/SASRec.pytorch.
HGN (Ma et al., 2019) propose a novel hierarchical gating mechanism to effectively capture both short-term and long-term user preferences in sequential recommendation tasks. Their model dynamically selects relevant interaction history at multiple temporal levels, improving next-item prediction accuracy. This approach outperforms state-of-the-art methods while maintaining efficiency and scalability for large-scale recommendation systems.
LightSANs (Fan et al., 2021) introduces a low-rank decomposition technique for self-attention networks, reducing their computational complexity while maintaining strong performance. This approach makes the model more efficient and scalable for large-scale recommendation tasks without compromising accuracy.
For the code implementation of Caser, HGN and LightSANs, we run the experiment based on the RecBole (Zhao et al., 2021)  [^10] https://github.com/RUCAIBox/RecBole/tree/master.
Self-supervised sequential recommendation methods
S3-Rec (Zhou et al., 2020) introduces self-supervised learning into a sequential recommendation by utilizing mutual information maximization (MIM) to learn better representations from user sequence data. The model incorporates four auxiliary self-supervised objectives: item cropping, item masking, item reordering, and segment prediction, to enhance the quality of learned item representations. By pre-training the model with these self-supervised tasks and then fine-tuning in the recommendation task, it achieves strong performance even with limited training data. The method demonstrates that self-supervised learning can effectively leverage the inherent supervisory signals within sequential data to improve recommendation quality.
DuoRec (Qiu et al., 2022) addresses the representation degeneration problem in sequential recommendation where item embeddings tend to be similar and occupy a narrow cone. Instead of traditional data-level augmentation (like masking or cropping), it introduces model-level augmentation using different Dropout masks to generate sequence representations. They also propose using sequences with the same target item as positive samples, which provides more meaningful contrasts. Through this contrastive regularization approach, DuoRec encourages a more uniform distribution of embeddings in the representation space, leading to better recommendation performance. The key innovation is tackling representation degeneration through model architecture rather than data augmentation while maintaining semantic consistency in the contrasting process.
MAERec (Ye et al., 2023) addresses limitations in sequential recommendation systems where other methods struggle with limited labels and noisy user behavior. Unlike previous approaches using manual contrastive learning strategies, MAERec implements a graph masked autoencoder framework that automatically identifies and focuses on meaningful item relationships. The model uses adaptive masking and task-specific regularization to ensure the learned representations align with recommendation goals while filtering out noise. By dynamically reconstructing masked item transitions rather than relying on hand-crafted data augmentation, MAERec achieves more robust performance across different recommendation scenarios without requiring manual heuristics. The approach demonstrates strong results in handling both data sparsity and noise while maintaining computational efficiency. The implementation of MAERec in PyTorch can be found at the URL  [^11] https://github.com/HKUDS/MAERec/tree/main.
For the code implementation of S3-Rec and DuoRec, we run the experiment based on the RecBole (Zhao et al., 2021)  [^12] https://github.com/RUCAIBox/RecBole/tree/master  [^13] https://github.com/RuihongQiu/DuoRec/tree/master.
@@
Open-P5 (Xu et al., 2023a) is an open-source platform introduced to catalyze research in LLM-based generative recommender systems. It supports key model architectures like T5 and Llama-2 across diverse public datasets, focusing on sequential and straightforward recommendation tasks. The platform emphasizes the role of item IDs through various indexing methods and offers a customizable, efficient, and standardized environment for developing and assessing recommender systems. The implementation of Open-P5 in PyTorch can be found at the URL   [^14] https://github.com/agiresearch/OpenP5.
E4SRec (Li et al., 2023a) integrate of Large Language Models (LLMs) into sequential recommendation systems, offering a significant leap in handling item IDs and personalization. In the original paper, they use Softmax layer to output each user-item prediction score. The implementation of E4SRec in PyTorch can be found at the URL  [^15] https://github.com/HestiaSky/E4SRec.
B.4 Model Efficiency
@@
To evaluate deployment efficiency in real-world scenarios, we compare the inference time between traditional recommendation approaches and LLM-based methods. Our experimental results demonstrate that SLMRec achieves comparable computational efficiency to traditional methods while delivering a substantial performance improvement of nearly 50% over SASRec.
@@
Table 8: Efficiency comparison of SASRec, MAERec and our SLMRec in terms of epoch-wise inference time (hours). These comparisons were conducted on a machine with an A100 GPU. During inference, models leverage parallel processing with a batch size of 512.
Method Inf time(h) Improv. (%)
SASRec 0.015 0.00
MAERec 0.061 11.96
SLMRec4←8 0.052 45.26
B.5 Ablation Study
We present the remaining ablation study results in Table 9. SLMRec, when enhanced with various knowledge regularizers (namely $\mathcal{D}_{cos},\mathcal{D}_{norm}$ and $\mathcal{L}_{ms}$), demonstrates improved performance. The regularizers $\mathcal{D}_{cos}$ and $\mathcal{D}_{norm}$ aid SLMRec in aligning its intermediate representations with those of the teacher model, thereby endowing it with more potent representational extraction capabilities. Meanwhile, $\mathcal{L}_{ms}$ steers the model to assimilate domain knowledge pertinent to recommendation systems within its preliminary layers.
Table 9: Experiment results (%) of ablation study.
SLMRec Music Sport
HR@1 HR@5 NDCG@5 MRR HR@1 HR@5 NDCG@5 MRR
+$\mathcal{D}_{cos}$ 5.62 8.78 7.23 7.81 6.25 9.25 7.76 8.41
+$\mathcal{D}_{cos}$,$\mathcal{D}_{norm}$ 5.95 9.26 7.65 8.23 6.61 9.82 8.24 8.87
+$\mathcal{D}_{cos}$,$\mathcal{L}_{ms}$ 5.69 8.94 7.36 7.91 6.51 9.39 7.96 8.62
+$\mathcal{D}_{cos}$,$\mathcal{D}_{norm}$,$\mathcal{L}_{ms}$ 5.72 9.15 7.48 8.03 6.62 9.83 8.25 8.89
Appendix C Extended Related Work
C.1 Depth-wise Knowledge of LLMs
The recent community interest stems from how linguistic properties and knowledge are encoded in language models.  (Meng et al., 2022; Dai et al., 2022; Jin et al., 2025) emphasize that knowledge localizes within the middle or final layers. On the other hand,  (Hase et al., 2024) attempts to perform knowledge editing and concludes that information may be stored non-locally across layers. What’s more,  (Men et al., 2024; Gromov et al., 2024) share a similar view that current pretraining methods are not properly leveraging the parameters in the deeper layers of the network or that the shallow layers play a critical role in storing knowledge. By contrast, we are the first to investigate which part of knowledge on the LLMs plays a key role, especially in the sequential recommendation scene.
C.2 Model Pruning
Model Pruning is a fundamental approach for reducing the size of a well-trained large model by removing unimportant parameters (Hassibi & Stork, 1992). Recent work has focused on applying pruning methods to the Transformer architecture (Vaswani et al., 2017). These works have studied different components of the model architecture for pruning, including dropping attention heads (Voita et al., 2019; Michel et al., 2019), dropping layers  (Fan et al., 2019; Zhang & He, 2020; Kim & Awadalla, 2020; Sajjad et al., 2023), dropping hidden xistates (Hou et al., 2020), replacing sparse weight matrices with smaller dense ones (Ashkboos et al., 2024), and combinations of these solutions. By contrast, our work performs layer removal through simple knowledge distillation, rather than more complex pruning techniques.
C.3 Parameter-Efficient Fine-Tuning (PEFT)
PEFT emerges as a novel technique for tailoring Large Language Models (LLMs) to specific tasks while ensuring minimal computational and memory costs (Houlsby et al., 2019; Lester et al., 2021; Hu et al., 2021b; Liu et al., 2022). In this work, we combine our method with the Low-Rank Adapters (LoRA) (Hu et al., 2021b) to reduce the memory and computation of the knowledge distillation process. Specifically, we freeze the pre-trained model and only tune a small set of additional trainable parameters.